OpenStudy (anonymous):

prove sin 90-theta as costheta

Parth (parthkohli):

\[\textbf{HINT}\]\[\sin(a-b) = \sin(a)\cos(b) - \cos(a)\sin(b) \]

OpenStudy (anonymous):

There are lots of ways to do this. Consider this drawing: